OpenStudy (anonymous):

which of these numbers are squared? how do you know? A.10 B.50 C.81 D.20

OpenStudy (anonymous):

They all might be squares, but only one is the square of a whole number.

OpenStudy (anonymous):

Think of what you would have to multiply to make those numbers.

OpenStudy (anonymous):

which one is it

OpenStudy (firejay5):

C. 81

OpenStudy (anonymous):

could you show me how you found ot

OpenStudy (anonymous):

Repeating: "Think of what you would have to multiply to make those numbers."

OpenStudy (anonymous):

ur in mr zimmers class!

OpenStudy (anonymous):

okay thanks guys

OpenStudy (firejay5):

9 x 9 = 81; when square rooting the same number has to be multiplied to get the answer. For Example: 4 * 4 = 16

OpenStudy (anonymous):

tthanx

OpenStudy (firejay5):

welcome

OpenStudy (anonymous):

these numbers are not square numbers. which 2 consucutive square numbers is each number between? describe your strategy? a) 12 b) 40 c) 75 d) 200

OpenStudy (anonymous):

help so confused

OpenStudy (anonymous):

ask mr zimmer

OpenStudy (anonymous):

im not

OpenStudy (anonymous):

This is for estimating square roots.

OpenStudy (anonymous):

For example, 12 is between 9 and 16. 9=3*#, 16=4*4, so the square-root of 12 is between 3 and 4.

OpenStudy (anonymous):

(oops, should be 9=3*3, not 3*#)

OpenStudy (anonymous):

sorry i still dont understand

OpenStudy (anonymous):

cliffsedge

OpenStudy (anonymous):

could you explain it

OpenStudy (anonymous):

OK, it's asking for perfect squares that are on either side of the given number. To make up my own example, let's say the number is 50. I know 50 isn't a perfect square because I can't think of any (whole) number multiplied by itself to make 50, but I can start guessing at perfect squares in that vicinity.

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

So let's see: 4*4=16, no that's way too low.. 10*10=100, so that's too big, so I know I'm looking for numbers between 4^2 and 10^2. 7*7=49 (that's pretty darn close), and 8*8=64. That's the best I can do, so I can say that 50 is between 49 and 64, which are consecutive square numbers.

OpenStudy (anonymous):

but then my textbook says what was your strategy whats the stragy to this

OpenStudy (anonymous):

I outlined my thought-process. Think about that and try a couple other examples yourself and then explain your own thought-process in your own words.

OpenStudy (anonymous):

you add 2 numbers togeather if one is close you put it there

OpenStudy (anonymous):

you multily 2 numbers by each other and if it close to the number you put on left side of the number and if it is highter right side

OpenStudy (anonymous):

Yes, you are looking for two consecutive numbers, like 4 and 5, 7 and 8, 20 and 21, etc. that when you square them (16 and 25, 49 and 64, 400 and 441 for my above examples), those squares will be on either side of the number you're working with.

OpenStudy (anonymous):

cliff sedge i gtg in like 10 mins could please give me thestregy you use

OpenStudy (anonymous):

I already did. I talked out my thought process and everything.

OpenStudy (anonymous):

i think its the squaring strategy